Using the above two equations to find the values of Qd, Qs, the market situation

(Shortage/Surplus/Equilibrium), and the Value of shortage or surplus if any, at the

following prices: 3.044, 3.65, 3.88, and 3.95.

the equations are QS=972+103.5p Qd=1722-141.5p


1
Expert's answer
2022-05-19T08:57:21-0400

"Q_s=972+103.5P,"

"Q_d=1722-141.5P,"

"Q_s=Q_d,"

"972+103.5P=1722-141.5P,"

"750=245P,"

"P=3.06,"

"Q=1288.8."